Fairfax Pizza is evaluating a project that would require an initial investment in equipment of 100,000 dollars and that is expected to last for 5 years. MACRS depreciation would be used where the depreciation rates in years 1, 2, 3, and 4 are 45 percent, 35 percent, 15 percent, and 5 percent, respectively. For each year of the project, Fairfax Pizza expects relevant, incremental annual revenue associated with the project to be 155,000 dollars and relevant, incremental annual costs associated with the project to be 135,000 dollars. The tax rate is 50 percent. What is (X plus Y) if X is the relevant operating cash flow (OCF) associated with the project expected in year 1 of the project and Y is the relevant OCF associated with the project expected in year 4 of the project?
Answer #1

X+Y = 32500+12500 = 45000

Year 1
Revenue 155,000
Less: Costs 135000
Depreciation=45%*100000 45000
Profit before tax -25,000
Tax savings -12500
Net Income -12,500
Add back:Depreciation 45000
OCF (X) 32,500
Year 4
Revenue 155,000
Less: Costs 135000
Depreciation=5%*100000 5000
Profit before tax 15,000
Tax savings 7500
Net Income 7,500
Add back:Depreciation 5000
OCF (Y) 12,500